# Dub-S2ST: Textless Speech-to-Speech Translation for Seamless Dubbing

### v1

@@ Abstract
This paper introduces a cross-lingual dubbing system that translates speech from one language to another while preserving key characteristics such as duration, speaker identity, and speaking speed. Despite the strong translation quality of existing speech translation approaches, they often overlook the transfer of speech patterns, leading to mismatches with source speech and limiting their suitability for dubbing applications. To address this, we propose a discrete diffusion-based speech-to-unit translation model with explicit duration control, enabling time-aligned translation. We then synthesize speech based on the predicted units and source identity with a conditional flow matching model. Additionally, we introduce a unit-based speed adaptation mechanism that guides the translation model to produce speech at a rate consistent with the source, without relying on any text. Extensive experiments demonstrate that our framework generates natural and fluent translations that align with the original speech’s duration and speaking pace, while achieving competitive translation performance.
@@ 1 Introduction
Recent advancements in translation systems and speech technologies have enabled numerous multimedia content to support multiple languages through automated speech dubbing. Cross-lingual dubbing Federico et al. (2020a); Wu et al. (2023), which replaces speech audio of one language with that of another, allows global audiences to consume content in their native languages. While this significantly reduces language barriers, ensuring effective dubbing requires meeting several specific criteria: maintaining duration, speaker identity, and speaking speed Brannon et al. (2023).
@@
Traditional dubbing systems typically employ a cascade of Automated Speech Recognition (ASR) Amodei et al. (2016); Baevski et al. (2020); Radford et al. (2023), Neural Machine Translation (NMT) Johnson et al. (2017); Stahlberg (2020); Fan et al. (2021); Costa-jussà et al. (2022), and Text-to-Speech (TTS) Wang et al. (2017); Ren et al. (2021); Wang et al. (2023); Tan et al. (2024b) modules. Although these cascaded systems demonstrate promising translation quality, they inherently lose critical speech-related information, such as speaker identity and prosody, due to the intermediate text representations Swiatkowski et al. (2023). Moreover, because text lacks precise duration information, these systems struggle to accurately match the duration and speaking pace of the original speech Sahipjohn et al. (2024). As a result, even with post-processing, the output often remains misaligned or unnatural, limiting their effectiveness for real-world dubbing Effendi et al. (2022).
@@
To address these limitations, Speech-to-Speech Translation (S2ST) systems Jia et al. (2019, 2022a); Seamless Communication (2025) have emerged and evolved into textless S2ST Lee et al. (2021); Li et al. (2023); Kim et al. (2024), aiming to translate speech directly without intermediate text. While recent approaches have achieved translation quality comparable to cascaded systems, most of them lack the ability to control speech duration during translation. As a result, the output typically requires post-processing like manual stretching or contracting to adjust the duration to the original speech for dubbing purposes. However, this process can degrade speech quality and lead to unnatural prosody and speaking pace.
@@
A key challenge underlying this issue lies in the limitations of existing training datasets. High-quality dubbing demands not only accurate translation but also faithful preservation of the source speech’s voice characteristics and speaking speed. However, it is inherently difficult to construct large-scale datasets containing the same speaker uttering aligned content in multiple languages. Consequently, most S2ST datasets rely on synthesized target speech Jia et al. (2022b) or web-crawled data Duquenne et al. (2023a); Seamless Communication (2025), which prioritize linguistic fidelity over speech-related information consistency. These datasets tend to exhibit discrepancies in speaker identity and speaking speed, making it challenging for models to learn how to preserve such attributes. Despite progress in translation accuracy, existing S2ST models trained on these datasets remain suboptimal for seamless dubbing.
@@
In response, we propose Dub-S2ST, a novel textless S2ST framework specifically designed for dubbing applications, effectively leveraging existing datasets. To mitigate the ambiguity caused by discrepancies between source and target speech, we eliminate variations in the target that deviate from the paired source. Specifically, we first convert continuous speech to discrete units that retain rich semantic features and minimal acoustic variations Lakhotia et al. (2021); Lee et al. (2021). We then apply our unit-based speed adaptation strategy to adjust the target’s speaking rate to the source. Using the processed data, we develop a speech-to-unit translation model trained with a discrete diffusion objective Austin et al. (2021). We also introduce Diffusion Transformer Peebles and Xie (2023) for this model, which allows to accurately predict speech units conditioning on diffusion timestep. Moreover, our model inherently supports duration control by using predetermined lengths based on the speech duration. Finally, we incorporate a conditional flow matching (CFM) Lipman et al. (2023); Mehta et al. (2024) based synthesizer that generates high-quality speech conditioned on the translated speech units and the original source speech, closely resembling the original speaker’s identity.
@@
Figure 1: Dub-S2ST framework. (a) avg. speed indicates average unit speed calculate in unit-based speed adaptation. (b) $h_{src}$ indicates the encoded source speech features from the encoder, and $t_{dec}$ is the timestep information used to train discrete-diffusion decoder. (c) $spk_{src},X_{src}$ represent speaker embedding and mel-spectrogram from the source speech as conditions, respectively. $t_{u2s}$ is the timestep information used to train unit-to-speech synthesizer.
@@ 2.1 Cross-lingual Dubbing
Dubbing is a post-production process in which the original spoken dialogue in multimedia content is replaced with speech in another language, while preserving the temporal alignment and naturalness of the original speech Orero (2004). Early automatic dubbing systems typically adopt cascaded S2ST architecture, combining ASR, NMT, and TTS Dureja and Gautam (2015). Maintaining such pipeline, recent efforts have enhanced each part of the components. Some approaches focus on improving prosodic alignment, aiming to synchronize the prosody of the generated speech with the original Federico et al. (2020b). Others leverage existing TTS models by modifying the duration module to generate speech that matches the original duration Effendi et al. (2022). Despite these advancements, relying on text as intermediate representation inherently limits temporal flexibility, highlighting the need for textless approaches that better preserve the naturalness of the source speech.
@@ 2.2 Speech-to-Speech Translation (S2ST)
Speech-to-Speech Translation (S2ST) aims to convert source speech into a target language while preserving linguistic content. Early systems adopted a cascaded architecture, integrating ASR, MT, and TTS modules Federico et al. (2020b); Lakew et al. (2022). To mitigate error propagation and latency issues inherent in cascaded systems, direct S2ST approaches have been introduced Jia et al. (2019, 2022a), demonstrating the feasibility of end-to-end speech translation. To further eliminate reliance on intermediate text, textless S2ST methods have emerged. S2UT Lee et al. (2021) proposes an autoregressive (AR) translation model that predicts deduplicated discrete speech units Lakhotia et al. (2021), which are then used to synthesize target speech. UTUT Kim et al. (2024) extends this framework to support many-to-many language translation. Subsequent works like TranSpeech Huang et al. (2023) explore non-autoregressive (NAR) models to achieve faster decoding compared to AR models, and DiffNorm Tan et al. (2024a) enhances TranSpeech by introducing techniques to normalize acoustic variations in the target speech. More recently, CTC-S2UT Fang et al. (2024) incorporates CTC Graves et al. (2006)-based unit reduction to improve translation performance. Unlike previous approaches, we leverage a NAR-based model without deduplicating speech units. This enables to preserve speech temporal information and explicitly control the output duration, making our framework naturally suitable for dubbing scenarios.
@@ 2.3 S2ST Datasets
Recent advancements in S2ST have been facilitated by the development of specialized datasets. VoxPopuli Wang et al. (2021) is a multilingual corpus containing aligned speech-to-speech pairs derived from simultaneous interpretation by human interpreters at European Parliament events. This dataset provides realistic translation pairs, however, it is limited in terms of domain coverage and scalability. A common approach to handle these issues is synthesizing target speech from translated text Jia et al. (2019, 2022a). For instance, CVSS dataset Jia et al. (2022b) is built by converting the text from the speech-to-text translated corpus CoVoST2 Jia et al. (2022b) into speech using a pretrained TTS model. Another approach is based on data mining. SeamlessAlign dataset Seamless Communication (2025) leverages web-crawled multimodal translation data and employs a unified speech-text similarity model Duquenne et al. (2023b) to effectively pair speech-speech segments, resulting in approximately 29,000 hours of paired S2ST data. However, these datasets still lack consistency in speech attributes between source and target. Our goal is to develop a framework that can generate dubbed speech that closely follows the original, even under such conditions.
@@ 3.1 Data Preparation
Speech Unit Extraction. The choice of target speech representation plays a critical role in determining the quality and accuracy of S2ST. While continuous features allow for straightforward generation, they are often susceptible to noise and mispronunciations, resulting in reduced intelligibility and naturalness. In contrast, discrete features leverage pretrained speech Self-Supervised Learning (SSL) models followed by quantization Hsu et al. (2021); Lee et al. (2021). Empowered by large-scale SSL training and discretization, this approach effectively captures phonetic information while minimizing speaker-dependent attributes such as timbre and pitch Lakhotia et al. (2021).
To focus explicitly on the linguistic components of speech during translation, the proposed method adopts units extracted using mHuBERT Lee et al. (2022), followed by k-means clustering for quantization. This cascaded unit extraction process facilitates robust representation of linguistic content while suppressing paralinguistic variations.
Unit-Based Speed Adaptation. Speaking speed is a essential factor in dubbing, as perceptual quality significantly degrades when the speed of translated speech deviates from the original Orero (2004). While signal processing techniques can be applied to adjust speech duration, they often compromise naturalness and intelligibility. Another strategy involves guiding models using syllable- or phoneme-based speed metrics, since those measures directly indicate speaking pace across langauges Seamless Communication (2025); however, such methods are inapplicable in textless systems.
To address these limitations, we propose a unit-based speed adaptation method that adjusts the repetition of speech units based on the speaking speed ratio between source and target speech. This method is inspired by unit deduplication (i.e., repetition removal) used in recent S2ST models Lee et al. (2021); Huang et al. (2023); Fang et al. (2024) to preserve synthesis quality while minimizing redundancy. We hypothesize that the reduced sequence $\hat{L}$ captures a distinct set of pronunciations, and the ratio against original length $L$ serves as an implicit estimate of speaking speed $r=\frac{\hat{L}}{L}$.
After extracting speech unit sequences from both source and target, we adjust the target sequence by applying the speed ratio $\frac{r_{src}}{r_{tgt}}$, modifying only the number of unit repetitions to align the target speed with the source. This adaptation preserves linguistic content while controlling speaking rate. Note that, as depicted in Figure 1, the speed adaptation method does not force the target speech units to have the same length as the source. It rather changes the rate of repeating units, so that it only alters the implicit speaking pace. Training the model on speed-aligned sequences enables it to generate translated speech that naturally and consistently mirrors the source speaking speed.
@@ 3.2 Speech-to-Unit Translation
Accurate dubbing requires the translated speech to match the duration of the source utterance. To fulfill this requirement, we propose a speech-to-unit translation model whose decoder is implemented as a discrete diffusion Austin et al. (2021) generator conditioned on the source speech. The decoder leverages Diffusion Transformer (DiT) Peebles and Xie (2023) layers that operate on variable-length sequences, with self-attention replaced by cross-attention to integrate source speech features throughout the generation. During inference, the decoder is initialized with all masked units with length identical to that of the source speech, and iteratively transforms it into speech units. Because the generation is explicitly conditioned on the source speech features, the output sequence faithfully represent the translated content while strictly preserving the original timing structure.
The training is conducted by first masking units based on a mask schedule $\gamma(t_{dec})$, where both $\gamma$ and $t_{dec}$ ranges from $(0,1)$. Each unit is independently masked with probability $\gamma(t_{dec})$ and remains unmasked with probability $1-\gamma(t_{dec})$, and the decoder predicts the original target units from the partially masked sequence. The decoder is optimized using the following cross-entropy loss:
@@
To enhance the decoder’s reliance on source speech representations, we initialize the encoder with weights from a pretrained autoregressive speech-to-unit translation model and fine-tune the entire weights. Additionally, we apply label-smoothing Szegedy et al. (2016) with a factor empirically set to 0.01 to improve generalization.
@@ 3.3 Unit-to-Speech Synthesis
A key factor in the perceptual quality of dubbing is the similarity between the translated speech and the original utterance. To convert semantic units to speech while preserving the acoustic characteristics of the source speech, we employ a unit-to-speech synthesizer based on Optimal Transport Conditional Flow Matching (OT-CFM). The synthesizer is implemented as a U-Net Rombach et al. (2022) architecture, where each layer is a block comprising Convoluational and Transformer layers. Downsampling and upsampling are performed in the latent space, allowing the model to efficiently reconstruct fine-grained temporal structure.
@@
The decoder is conditioned on a tuple of features:
$\boldsymbol{c}=\mathrm{concat}(spk_{\text{src}},X_{\text{src}},\mathrm{unit}_{\text{tgt}})$ (5)
where $\mathrm{unit}_{\text{tgt}}$ is the generated target units from the speech-to-unit module, $spk_{\text{src}}$ is the speaker embedding of the source speech, and $X_{\text{src}}$ is its mel-spectrogram. These features are concatenated to enable in-context learning: the sampled prior is transformed into a mel-spectrogram conditioned on both speaker identity and prosodic information from the reference utterance.
Although directly training the module on S2ST data is possible, such datasets often contain noise and reverberation, which can impair synthesis quality. To address this, we initialize the model with weights trained on multilingual corpus, and fine-tune it with our own semantic units. This approach allows for robust zero-shot synthesis across diverse speakers and languages. The generated mel-spectrogram is then converted to audible waveform via pretrained HiFi-GAN Kong et al. (2020).

### v2

@@ Abstract
This paper introduces a cross-lingual dubbing system that translates speech from one language to another while preserving key characteristics such as duration, speaker identity, and speaking speed. Despite the strong translation quality of existing speech translation approaches, they often overlook the transfer of speech patterns, leading to mismatches with source speech and limiting their suitability for dubbing applications. To address this, we propose a discrete diffusion-based speech-to-unit translation model with explicit duration control, enabling time-aligned translation. We then synthesize speech based on the translated units and source speaker’s identity using a conditional flow matching model. Additionally, we introduce a unit-based speed adaptation mechanism that guides the translation model to produce speech at a rate consistent with the source, without relying on any text. Extensive experiments demonstrate that our framework generates natural and fluent translations that align with the original speech’s duration and speaking pace, while achieving competitive translation performance. The code is available at https://github.com/kaistmm/Dub-S2ST.
@@ 1 Introduction
Recent advancements in translation systems and speech technologies have enabled a vast amount of multimedia content to support multiple languages through automated speech dubbing. Cross-lingual dubbing federico2020speech; wu2023videodubber, which replaces speech audio of one language with that of another, allows global audiences to consume content in their native languages. While this significantly reduces language barriers, ensuring effective dubbing requires meeting several specific criteria: maintaining duration, speaker identity, and speaking speed brannon2023dubbing.
@@
Traditional dubbing systems typically employ a cascade of Automated Speech Recognition (ASR) amodei2016deep; baevski2020wav2vec; radford2023robust, Neural Machine Translation (NMT) johnson2017google; stahlberg2020neural; fan2021beyond; costa2022no, and Text-to-Speech (TTS) wang2017tacotron; ren2021fastspeech; wang2023neural; tan2024naturalspeech modules. Although these cascaded systems demonstrate promising translation quality, they inherently lose critical speech-related information, such as speaker identity and prosody, due to the intermediate text representations swiatkowski2023cross. Moreover, because text lacks precise duration information, these systems struggle to accurately match the duration and speaking pace of the original speech sahipjohn2024dubwise. As a result, even with post-processing, the output often remains misaligned or unnatural, limiting their effectiveness for real-world dubbing effendi2022duration.
@@
To address these limitations, Speech-to-Speech Translation (S2ST) systems jia2019translatotron; jia2022translatotron; barrault2025seamless have emerged and evolved into textless S2ST lee2021textless; li2023textless; kim2024textless, aiming to translate speech directly without intermediate text. While recent approaches have achieved translation quality comparable to cascaded systems, most of them lack the capability to control speech duration during translation. As a result, the output typically requires post-processing like manual stretching or contracting to match the original speech duration for dubbing purposes. However, this process can degrade speech quality and lead to unnatural prosody and speaking pace.
@@
A key challenge underlying this issue lies in the limitations of existing training datasets. High-quality dubbing demands not only accurate translation but also faithful preservation of the source speech’s voice characteristics and speaking speed. However, it is inherently difficult to construct large-scale datasets containing the same speaker uttering aligned content in multiple languages. Consequently, most S2ST datasets rely on synthesized target speech jia2022cvss or web-crawled data duquenne2022speechmatrix; barrault2025seamless, which prioritize linguistic fidelity over speech-related information consistency. These datasets tend to exhibit discrepancies in speaker identity and speaking speed, making it challenging for models to learn how to preserve such attributes. Despite progress in translation accuracy, existing S2ST models trained on these datasets remain suboptimal for seamless dubbing.
@@
In response, we propose Dub-S2ST, a novel textless S2ST framework specifically designed for dubbing applications, effectively leveraging existing datasets. To mitigate the ambiguity caused by discrepancies between source and target speech, we eliminate variations in the target that deviate from the paired source. Specifically, we first convert continuous speech to discrete units that retain rich semantic features and minimal acoustic variations lakhotia2021generative; lee2021textless. We then apply our unit-based speed adaptation strategy to adjust the target’s speaking rate to the source. Using the processed data, we develop a speech-to-unit translation model trained with a discrete diffusion objective austin2021structured. We also incorporate Diffusion Transformer peebles2023scalable, which allows the model to accurately predict speech units conditioned on diffusion timestep. Moreover, our model inherently supports duration control by using predetermined lengths based on the duration of source speech. Finally, we incorporate a conditional flow matching (CFM) lipman2023flow; mehta2024matcha-based synthesizer that generates high-quality speech conditioned on the translated speech units and the original source speech, closely resembling the original speaker’s identity.
@@
Figure 1: Dub-S2ST framework. (a) avg. speed indicates average unit speed calculated in unit-based speed adaptation. (b) $h_{src}$ indicates the encoded source speech features from the encoder, and $t_{dec}$ is the timestep information used to train discrete-diffusion decoder. (c) $spk_{src},X_{src}$ represent speaker embedding and mel-spectrogram from the source speech as conditions, respectively. $t_{u2s}$ is the timestep information used to train unit-to-speech synthesizer.
@@ 2.1 Cross-lingual Dubbing
Dubbing is a post-production process in which the original spoken dialogue in multimedia content is replaced with speech in another language, while preserving the temporal alignment and naturalness of the original speech orero2004topics. Early automatic dubbing systems typically adopt cascaded S2ST architecture, combining ASR, NMT, and TTS dureja2015speech. While maintaining this cascaded pipeline, recent efforts have focused on enhancing each component. Some approaches focus on improving prosodic alignment, aiming to synchronize the prosody of the generated speech with the original federico2020evaluating. Others leverage existing TTS models by modifying the duration module to generate speech that matches the original duration effendi2022duration. Despite these advancements, relying on text as intermediate representation inherently limits temporal flexibility, highlighting the need for textless approaches that better preserve the naturalness of the source speech.
@@ 2.2 Speech-to-Speech Translation (S2ST)
Speech-to-Speech Translation (S2ST) aims to convert source speech into a target language while preserving linguistic content. Early systems adopted a cascaded architecture, integrating ASR, MT, and TTS modules federico2020evaluating; lakew2022isometric. To mitigate error propagation and latency issues inherent in cascaded systems, direct S2ST approaches have been introduced jia2019translatotron; jia2022translatotron, demonstrating the feasibility of end-to-end speech translation. To further eliminate reliance on intermediate text, textless S2ST methods have emerged. S2UT lee2021textless proposes an autoregressive (AR) translation model that predicts deduplicated discrete speech units lakhotia2021generative, which are then used to synthesize target speech. UTUT kim2024textless extends this framework to support many-to-many language translation. Subsequent works like TranSpeech huang2023transpeech explore non-autoregressive (NAR) models to achieve faster decoding compared to AR models, and DiffNorm tan2024diffnorm enhances TranSpeech by introducing techniques to normalize acoustic variations in the target speech. More recently, CTC-S2UT fang2024ctc incorporates CTC graves2006connectionist-based unit reduction to improve translation performance. Unlike previous approaches, we leverage a NAR-based model without deduplicating speech units. This design preserves temporal information and enables explicit control over output duration, making our framework naturally suitable for dubbing scenarios.
@@ 2.3 S2ST Datasets
Recent advancements in S2ST have been facilitated by the development of specialized datasets. VoxPopuli wang2021voxpopuli is a multilingual corpus containing aligned speech pairs derived from simultaneous interpretation by human interpreters at European Parliament events. This dataset provides realistic translation pairs, however, it is limited in terms of domain coverage and scalability. A common approach to handle these issues is synthesizing target speech from translated text jia2019translatotron; jia2022translatotron. For instance, CVSS jia2022cvss is built by converting the text from the speech-to-text translation corpus CoVoST2 jia2022cvss into speech using a pretrained TTS model. Another approach is based on data mining. SeamlessAlign barrault2025seamless leverages web-crawled multimodal translation data and employs a unified speech-text similarity model duquenne2023sonar to effectively pair speech segments, resulting in approximately 29,000 hours of S2ST data. However, these datasets still lack consistency in speech attributes between source and target. Our goal is to develop a framework that can generate dubbed speech that closely follows the source, even under such conditions.
@@ 3.1 Data Preparation
Speech Unit Extraction. The choice of target speech representation plays a critical role in determining the quality and accuracy of S2ST. While continuous features allow for straightforward generation, they are often susceptible to noise and mispronunciations, resulting in reduced intelligibility and naturalness. In contrast, discrete features leverage pretrained speech Self-Supervised Learning (SSL) models followed by quantization hsu2021hubert; lee2021textless. Empowered by large-scale SSL training and discretization, this approach effectively captures phonetic information while minimizing speaker-dependent attributes such as timbre and pitch lakhotia2021generative.
To focus explicitly on the linguistic components of speech during translation, the proposed method adopts units extracted using mHuBERT lee2022direct, followed by k-means clustering for quantization. This cascaded unit extraction process facilitates robust representation of linguistic content while suppressing paralinguistic variations.
Unit-Based Speed Adaptation. Speaking speed is an essential factor in dubbing, as perceptual quality significantly degrades when the speed of translated speech deviates from the original orero2004topics. While signal processing techniques can be applied to adjust speech duration, they often compromise naturalness and intelligibility. Another strategy is to guide models using syllable- or phoneme-based speed metrics, since these provide an indication of speaking pace across languages barrault2025seamless; however, such methods are inapplicable in textless systems.
To address these limitations, we propose a unit-based speed adaptation method that adjusts the repetition of speech units based on the speaking speed ratio between source and target speech. This method is inspired by unit deduplication (i.e., repetition removal) used in recent S2ST models lee2021textless; huang2023transpeech; fang2024ctc to preserve synthesis quality while minimizing redundancy. We hypothesize that the reduced sequence $\hat{L}$ captures a distinct set of pronunciations, and the ratio against original length $L$ serves as an implicit estimate of speaking speed $r=\frac{\hat{L}}{L}$.
After extracting speech unit sequences from both source and target, we adjust the target sequence by applying the speed ratio $\frac{r_{src}}{r_{tgt}}$, modifying the number of unit repetitions to align the target speed with the source. The speed is normalized by the average speed of each language to mitigate cross-linguistic differences. This adaptation preserves linguistic content while controlling speaking rate. Note that, as depicted in Figure 1, the speed adaptation method does not force the target speech units to have the same length as the source. It rather changes the rate of repeating units, so that it only alters the implicit speaking pace. Training the model on speed-aligned sequences enables it to generate translated speech that naturally and consistently mirrors the source speaking speed.
@@ 3.2 Speech-to-Unit Translation
Accurate dubbing requires the translated speech to match the duration of the source utterance. To fulfill this requirement, we propose a speech-to-unit translation model whose decoder is implemented as a discrete diffusion austin2021structured generator conditioned on the source speech. The decoder leverages Diffusion Transformer (DiT) peebles2023scalable layers that operate on variable-length sequences, with additional cross-attention to integrate source speech features throughout the generation. During inference, the decoder takes a fully masked unit sequence with length identical to that of the source speech as its initial input and iteratively transforms it into speech units.
The training is conducted by first masking units based on a mask schedule $\gamma(t_{dec})$, where both $t_{dec}$ and $\gamma(t_{dec})$ range from $(0,1)$. Each unit is independently masked with probability $\gamma(t_{dec})$ and remains unmasked with probability $1-\gamma(t_{dec})$, and the decoder predicts the original target units from the partially masked sequence. The decoder is optimized using the following cross-entropy loss:
@@
To enhance the decoder’s reliance on source speech representations, we initialize the encoder from a pretrained autoregressive speech-to-unit translation model lee2022direct and fine-tune the entire weights. Additionally, we apply label-smoothing szegedy2016rethinking with a factor empirically set to 0.01 to improve generalization.
@@ 3.3 Unit-to-Speech Synthesis
A key factor in the perceptual quality of dubbing is the similarity between the translated speech and the original utterance. To convert semantic units to speech while preserving the acoustic characteristics of the source speech, we employ a unit-to-speech synthesizer based on Optimal Transport Conditional Flow Matching (OT-CFM). The synthesizer is implemented as a U-Net rombach2022high architecture, where each layer is a block comprising Convolutional and Transformer layers. Downsampling and upsampling are performed in the latent space, allowing the model to efficiently reconstruct fine-grained temporal structure.
@@
The condition $\boldsymbol{c}$ consists of the unit embedding sequence from source and target speech, the source speaker embedding $spk_{\text{src}}$ extracted with a pretrained speaker verification model wang2023cam++, and the source mel-spectrogram $X_{\text{src}}$. The mel-spectrogram is padded to match the length of the unit sequence, while the speaker embedding is repeated accordingly. These features are concatenated channel-wise to enable in-context learning: the sampled prior is transformed into a mel-spectrogram conditioned on both speaker identity and prosodic information from the source speech.
Although directly training the module on S2ST data is possible, such datasets often contain noise and reverberation, which can impair synthesis quality. To address this, we initialize the model from a TTS model du2024cosyvoice trained on multilingual corpus, and fine-tune it with necessary adaptations for our semantic unit input. This approach allows for robust zero-shot synthesis across diverse speakers and languages. The generated mel-spectrogram is then converted to audible waveform via a pretrained HiFi-GAN kong2020hifi.